OpenStudy (hang254):

The position of an object at time t is given by s(t) = -2 - 6t. Find the instantaneous velocity at t = 2 by finding the derivative.

OpenStudy (hang254):

@Directrix

Directrix (directrix):

Do you know the derivative of this: s(t) = -2 - 6t

Directrix (directrix):

@hang254 Yes or No

OpenStudy (hang254):

-6 ?

OpenStudy (hang254):

@SolomonZelman

OpenStudy (hang254):

@micahm

Directrix (directrix):

s(t) = -2 - 6t s'(t) = -6 What is s'(2) = @hang254